OpenStudy (anonymous):

Solve 5x – 3 = –3x + 5

OpenStudy (anonymous):

do you have any idea how to do this one?

OpenStudy (anonymous):

well i'll go ahead and tell you the direction to go:

OpenStudy (anonymous):

by finding the value of X

OpenStudy (anonymous):

yes good. we need to isolate it, can you tell me how you would do that? go about it the easiest way possible

OpenStudy (anonymous):

no i dont know how?

OpenStudy (anonymous):

ok, we would take the -3 on the left side and add it to both side of the equation so we get 5x - 3 + 3 = -3x + 5 +3 so that gives up 5x = -3x + 5 + 3

OpenStudy (anonymous):

or just 5x = -3x + 8

OpenStudy (anonymous):

can you do it from here? what do you think you would do next

OpenStudy (anonymous):

okay i understand, whats next?

OpenStudy (anonymous):

alright great. so then we still need to keep isolating that x, we aren't done yet. we will add 3x to both sides and get 3x + 5x = -3x + 8 + 3x and that will give us 8x = 8

OpenStudy (anonymous):

X=1

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ThankYou so much!

OpenStudy (anonymous):

if you want to check yoruself you can put 1 where the x is in the equation and solve it that way and you should get the same number on both sides

OpenStudy (anonymous):

you are very welcome, have a good night